If 44.5 mol of an ideal gas occupies 77.5 L at 61.00 °C, what is the pressure of the gas? Number atm
Answer #1

PV = nRT

where, P = pressure = ?

V = volume = 77.5 L

n = number of moles = 44.5 mol

R = Gas constant

T = temperature = 61.00 + 273 = 334 K

P * 77.5 = 44.5 * 0.0821*334

77.5 * P = 1220

P = 1220 / 77.5 = 15.7 atm

Therefore, pressure = 15.7 atm